OpenStudy (anonymous):

does x^2-x-12=(x-3)(x+4)?

OpenStudy (angela210793):

(x-3)(x+4)=x^2-4x-3x+12=x^2-7x-12 so no

OpenStudy (anonymous):

scarlett u have alredy asked it 3 times i think u can solve it now :)

OpenStudy (anonymous):

(x-3)(x+4)= x^2 +x -12 (So no)